Hi, can I have some help with this problem? Step by Step explanations would be very helpful!! Thanks

A wire having a linear mass density of 1.00 g/cm i
Answer #1

The wire is experiencing a force in the north direction .The force experienced by a current carrying conductor placed in a magnetic field is given as
F = B i l
The direction of the current and the direction of magnetic field is perpendicular , so we dropped the cross product
Friction acts opposite to the direction of movement
The magnitude of friction is f = \muk x N
Where  \muk is the coefficient of the kinetic friction and N is the normal force
To find the minimum magnetic field we equate these two forces
B i l =   \muk x m x g
B =   \muk x m x g / i l
\muk = 0.200 , linear mass density m / l = 1 g /cm3  , i = 1.5 A ,g = 9.8 m /s2
B = 0.2 x 1000 x 9.8 / 1.5 = 1306.6 T
b) The direction of the magnetic field should be into the surface